OpenStudy (anonymous):

Jennifer belongs to a gym that requires a monthly membership fee of $60 plus an additional $15 fee for each yoga class she attends. Which of the following slope-intercept form equations models the total amount that Jennifer pays monthly?

OpenStudy (anonymous):

A. y = –60x + 15 B. y = –15x + 60 C. y = 15x + 60 D. y = 60x + 15

OpenStudy (anonymous):

same format as last question but this time: $60 will be your one time fee. $15 will be per each yoga class. let y=monthly cost let x=number of yoga classes taken. y=15x+60 choice C

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no prob!